1/ x-1 + 1/2 = 2/ x^2 -1
I am trying to help my daughter with this problem, but I can't get it to work out. Can you please show how to explain. Thank you.

Found 2 solutions by rwm, unlockmath:
Answer by rwm(914)

Enclose the numerators and denominators in parentheses and then resubmit
or send it to me

Answer by unlockmath(1688)

Hello,
For this one: 1/ x-1 + 1/2 = 2/ x^2 -1 I'm assuming the x-1 and the x^2-1 are in the denominator. If so we can do the following:
Multiply by (x-1)(x+1) to get:
(x+1)+ 1/2(x-1)(x+1)
NOw let's multiply by 2 to get rid of the fraction:
2(x+1)+(x-1)(x+1) = 4
Expand this out:
x^2+2x+1=4
Now subtract 4 from both sides to get:
x^2+2x-3=0
Now factor:
(x-1)(x+3)=0
Solve for x
x=1
x=-3
x=1 doesn't work but
x=-3 works so there we go.
Make sense?
RJ
www.math-unlock.com